OpenStudy (anonymous):

2sinxcscx-cscx=4sinx-2.. find the solutions of the equation that are in the interval [0, 2pi)

OpenStudy (dumbcow):

\[2\sin x \frac{1}{\sin x} - \frac{1}{\sin x} = 4\sin x -2\] \[2-\frac{1}{\sin x} = 4\sin x -2\] \[-\frac{1}{\sin x} = 4\sin x -4\] \[-1 = 4\sin^{2} x -4\sin x\] \[4\sin^{2} x -4\sin x +1 = 0\] factor \[(2\sin x -1)^{2} = 0\] \[2\sin x = 1\] \[\sin x = \frac{1}{2}\] \[\rightarrow x = \frac{\pi}{6}, \frac{5\pi}{6}\]